dA run progress

From: Flemming Videbaek (videbaek@rcf.rhic.bnl.gov)
Date: Sat Mar 01 2003 - 14:43:24 EST

  • Next message: Apache: "Shift report 20030301 08:00-16:00"
    Dear Collaborator
    
    I hope you are following the run-progress via the shift reports and the
    e-logs. Let me summarize briefly where we are in the program.
    
    Most of soft physics program has been completed. The exception is mostly
    that we decided due to the problems with a missing TD1 slat in the FFS
    trigger to repeat the 8 deg runs. These are currently underway, and with
    the
    present most improved stores (ZDC~9K initial and reasonable backgrounds)
    these run can be expected to be completed in a few (2-3) good stores.
    
    Then the higher pt program can start i.e. the 4 and 3 degree higher field
    (1/4) 1/2 and either 2/3 or 1/1 settings depending on the actual rates
    observed can is up. The 12 deg higher field setting will be done
    subsequently; if both polarities can be done depends on RHIC progress and
    on possible decisions to extend the dA run. Rate estimates indicates that
    about a week is needed to do one pol at 12 deg.
    For the mid-rapidity region we have alreay a quite good sample ~3M triggers
    (~.8M tracks) at y~0. With the C4 Cherenkov detector physical installed on
    the platform (with 1/2 the PMTs so far) we can hope to have it operational
    sometime this week and can pursue measurements of high pt pions using both
    C4 and the tof extension also in place.
    
    We continue to have experimental problems of various kinds that have
    impacted the amount of data we have taken. Lots of this has been in regard
    to HV problems with DCs and TPCs both related to detector, HV supplies and
    to misc. DAQ hickups.
    
    To improve the experiment uptime I urge the shifters to continue to
    
    -- beware of present detector status and problems by reading e-logs and
    discuss at start of shift in detail with previous shift status of
    experiment an experimental program. Please do also carefull monitor the
    progress checking the monitor plots and printouts, as well as pying
    carefull attention to beam quality (backgrounds) and
    
    -- In case of problems make sure to consult the 'shift guide' and various
    notes poste in the CH. The shift guide 'problem section' has been updated
    today and includes description of problems and solutions to misc. problems
    that we have had recently.
    
    regards
       Flemming
    
    
    
    -- 
    Flemming Videbaek
    Brookhaven National Laboratory
    Physics Department
    tlf 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    


    This archive was generated by hypermail 2.1.5 : Sat Mar 01 2003 - 14:44:02 EST